OpenStudy (anonymous):

how many grams of CO2 gas of 40 degree celcius does it take to fill a 1.0 liters flask to a pressure of 1.5 atm?

OpenStudy (chmvijay):

you can use gas equation PV=nRT put the values of P-1.5 atm ,v-1 lit ,T-40 °C and R gas constant value and find n which gives you no of moles of CO2 then moles you multiply by molecular weight uou get amount of CO2 in grams

OpenStudy (chmvijay):

did you get it

OpenStudy (anonymous):

...could u solve for it so that i can understand..please :DD

OpenStudy (anonymous):

i try to solve.. i get 890.50 is that right??

OpenStudy (anonymous):

how it will become??

OpenStudy (anonymous):

how did u get the moles?

OpenStudy (chmvijay):

the n value in the above equation will be alwas in mole unit only

OpenStudy (chmvijay):

solve for n in the gas equation

OpenStudy (anonymous):

please give me the solution on how you do the gas equation..please..

OpenStudy (chmvijay):

hey sory i was wrong it will come 0.059 moles and in grams as 2.59 gram sory i made mistake in calculation thank you for that

OpenStudy (anonymous):

...how did u get that moles??

OpenStudy (chmvijay):

PV=nRT where n is noumber of moles of gas 1.5*1litre =n *0.082*313 n=(1.5*1) / (0.082*313) =1.5/25.66 n =0.059 this mean 0.059 moles of gas

OpenStudy (anonymous):

@chmvijay thank u very much!! may the god bless u!! can u please get it in grams??

OpenStudy (chmvijay):

you multiply that with its molecular weigh you will get it 44*0.059=2.59 gram

OpenStudy (anonymous):

thanks!!